You have just committed to a building project where you would need to make the

following payments to your contractor in the next 10 years:

a. An immediate down payment of $200,000.

b. $130,000 at the end of year 3.

c. $96,000 at the end of year 5.

d. $312,000 at the end of year 8.

e. $1,250,000 at the end of year 10.

Savers bank has offered you a bank account which will pay you 7.25% compounded

weekly. How much money do you need to deposit into this account today, so you can

make the above payments over the next 10 years?
Answer #1

Given data, the savers bankers offered interest rate @7.25percent per annum compounding weekly

Weekly interest rate = 7.25/52 = 0.14 percent

Formula: Future value = present value (1+r)n

By applying above formula:

1250000 = P.V(1+0.0014)520

Present value = 6,03,900 ( By rounding off)  

312000 = P.V(1+0.0014)416

Present value = 1,74,341

96000 = P.V(1+0.0014)260

Present value= 66,727

130000 = P.V (1+0.0014)156

Present value = 77,178

Down payment = 2,00,000

Total amount to be deposited = $ 11,22,146.00

Notes to sum:

Here n = weeks×number of years

1) n = 52×10 = 520

2)n = 52×8 = 416

3)n = 52×5 = 260

4)n = 52×3 = 156

​​​​​​

​​​​​​